In a party for 30 young children, 12 were served Coca-Cola drink, while 14 were served pepsi-cola . If 9 children were served Coca-Cola but not with pepsi-cola , find
1)the number of children served both drinks
2).the number of children served at least one type of drink
3).the number of children not served any of the drink

Answer by ikleyn(52781)

.
In a party for 30 young children, 12 were served Coca-Cola drink, while 14 were served pepsi-cola .
If 9 children were served Coca-Cola but not with pepsi-cola , find
1) the number of children served both drinks
2) the number of children served at least one type of drink
3) the number of children not served any of the drink
~~~~~~~~~~~~~~


(1)  the number of children served both drinks                = 12 - 9 = 3.          ANSWER


(2)  the number of children served at least one type of drink = 12 + 14 - 3 = 23.    ANSWER


(3)  the number of children not served any of the drink       = 30 - 23 = 7.         ANSWER

Solved.     //     All questions are answered.

------------------

To see other similar problems solved and to make your horizon wider, look into the lesson
    - Counting elements in sub-sets of a given finite set
in this site.